# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v1

@@ 3.2.1 Global convergence
    if $x^{k}\rightarrow x^{*}$, $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$, then $x^{*}$ is a $K$-stationary point to (VOP).
@@
The Assumption 3.1(i) is a standard condition for nonconvex cases. Moreover, the Assumption 3.1(ii) holds for many descent methods, such as $K$-steepest descent method, Newton method.
@@
$-\infty<\max_{c^{*}\in C}\left\langle{c^{*},F^{*}-F(x^{0})}\right\rangle\leq\max_{c^{*}\in C}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\sum_{k=0}^{\infty}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle.$
This, together with the fact that $G_{k}(x^{k+1})\preceq_{K}G_{k}(x^{k})=0$, implies $\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$. For the accumulation point $x^{*}$, there exists an infinite index set $\mathcal{K}$ such that $x^{k}\stackrel{{\scriptstyle\mathcal{K}}}{{\longrightarrow}}x^{*}$. Therefore, by Assumption 3.1(ii), we conclude that $x^{*}$ is a $K$-stationary point. ∎
@@ 3.2.2 Strong convergence
Suppose that Assumption 3.1 holds and $F(\cdot)$ is $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\bm{\ell}\preceq_{K}\bm{\mu}$. Then, the following statements hold:
@@
    $u_{0}(x^{k})\leq\frac{\ell_{\max}R^{2}}{2k},~{}\forall k\geq 1$, where $\ell_{\max}:=\max_{c^{*}\in C}\left\langle{c^{*},\bm{\ell}}\right\rangle$, $R:=\{\left\lVert{x-y}\right\rVert:x,y\in\mathcal{L}_{F}(x^{0})\}$, and
    $u_{0}(x^{k}):=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$
@@
Furthermore, we use the fact that $\bm{\ell}\preceq_{K}\bm{\mu}$ to get
@@
(ii) Since $\bm{\ell}\preceq_{K}\bm{\mu}$, we use inequality (3) and the definition of $\ell_{\max}$ to obtain
@@
Denote $\hat{c}^{*}_{k}:={\sum_{s=0}^{k-1}c^{*}_{s}}/{k}$, it follows by the convexity of $C$ that $\hat{c}^{*}_{k}\in C$. Therefore, we conclude that
@@
Select $y^{k}\in\mathop{\arg\max}_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$, it holds that
$u_{0}(x^{k})$ $=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle=\min_{c^{*}\in C}\left\langle{c^{*},F(x^{k})-F(y^{k})}\right\rangle$
@@ 3.2.3 Linear convergence
Suppose that Assumption 3.1(ii) holds $F(\cdot)$ is strongly $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\bm{\ell}\prec_{K}\bm{\mu}$. Then, the following statements hold:
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{\max\limits_{c^{*}\in C}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
@@
4 Descent method for VOPs without line search
@@ 4.1 $K$-steepest descent method for VOPs without line search
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}.$ (4)
Select a vector $e\in{\rm int}(K)$, and denote $C_{e}=\{c^{*}\in K^{*}:\left\langle{c^{*},e}\right\rangle=1\}$. If we set $C=C_{e}$ in (4), then the $K$-steepest descent direction can be reformulated as the optimal solution of
@@
Since $F(\cdot)$ is strongly $K$-convex, $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$, and Assumption 3.1 holds in this case, using $C=C_{e}$, then Theorem 3.3 (i) and (ii) reduce to the assertions (i) and (ii), respectively. ∎
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$ a base of +$\mathbb{R}^{m}_{+}$, the convergence rate result in Lemma 4.1 reduces to that of [39, Theorem 5.3] with $g(\cdot)=0$. It is worth noting that the rate of convergence is related to $e$, to the best of our knowledge, apart from $e=(1,...,1)^{T}$, it remains an open problem for the better choice of the parameter in MOPs.
@@ 4.3 Equivalence between Algorithms 2 and 3
If $K=\mathbb{R}^{m}_{+}$, and $\Delta_{m}^{\bm{\ell}}:=\{c^{*}\in\mathbb{R}^{m}_{++}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}$, the Algorithm 4 reduces to [8, Algorithm 5] with $g(\cdot)=0$. Interestingly, the relations between Algorithms 2, 3 and 4 only depends on the choice of $e$. If $e=\bm{\ell}$, Algorithms 2, 3 and 4 are equivalent. On the other hand, for the open problem of the better choice of $e$ in Remark 4.2, we give a partially positive answer with $e=\bm{\ell}$.
@@
Even though Algorithms 3 and 4 enjoy similar improved linear convergence, the computational cost of subproblems in Algorithms 4 is often cheaper.
5 Descent method for VOPs with line search
In the preceding section, the majorization-minimization optimization methods are devised based on majorizing surrogates, which might be too conservative due to the global upper bounds. From a majorization-minimization perspective, choosing a non-majorizing surrogate function may provide better performance.
@@ 7             $x^{k+1}:=x^{k}+t_{k}d^{k}$
$G_{k,e/t_{k}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2t_{k}}\left\lVert{x-x^{k}}\right\rVert^{2}e.$ (9)
We can show that $G_{k,e/t_{k}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
@@
Let $G_{k,e/t_{k}}(\cdot)$ be defined as (9). Then, the following statements hold.
@@
    If $F(\cdot)$ is $K$-convex, then $G_{k,e/t_{k}}(\cdot)\in\mathcal{S}_{e/t_{k},e/t_{k}}(F,x^{k})$.
@@
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,e/t_{k}}(\cdot)\in\mathcal{S}_{e/t_{k}-\bm{\mu},e/t_{k}}(F,x^{k})$.
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$, the convergence rate result in Lemma 5.1 reduces to that of [18, Theorem 4.2].
Proof.
@@ 8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
The stepsize has the following lower bound.
@@
The stepsize generated in Algorithm 6 satisfies $t_{k}\geq t_{\min}:=\min\left\{\min\limits_{c^{*}\in C_{e}}\frac{\gamma\left\langle{c^{*},e_{k}}\right\rangle}{\left\langle{c^{*},\bm{\ell}}\right\rangle},1\right\}$.
@@
$G_{k,e_{k}/t_{k}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2t_{k}}\left\lVert{x-x^{k}}\right\rVert^{2}e_{k}.$ (10)
@@
Let $G_{k,e_{k}/t_{k}}(\cdot)$ be defined as (10). Then, the following statements hold.
@@
    If $F(\cdot)$ is $K$-convex, then $G_{k,e_{k}/t_{k}}(\cdot)\in\mathcal{S}_{e_{k}/t_{k},e_{k}/t_{k}}(F,x^{k})$.
@@
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,e_{k}/t_{k}}(\cdot)\in\mathcal{S}_{e_{k}/t_{k}-\bm{\mu},e_{k}/t_{k}}(F,x^{k})$.
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t_{k}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
@@
By set $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$, we can derive that $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-{1}/{\kappa_{F,\preceq_{K}}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$. Therefore, convergence rate in Lemma 5.2(ii) reduces to that of Lemma 4.3(ii). Intuitively, to explore the local curvature information of $F(\cdot)$, we can devise a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$ with $\bm{\mu}\preceq_{K}e_{k}/t_{k}\preceq_{K}\bm{\ell}$. In this case, the linear convergence rate of Algorithm 6 can be further improved by using a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$.
@@
6 Descent method for VOPs with polyhedral cones

### v2

@@ 3.2.1 Global convergence
    There exist two bases $C$ and $\tilde{C}$ such that
    $\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle$
    holds for all $k\geq 0$. If $x^{k}\rightarrow x^{*}$, $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$, then $x^{*}$ is a $K$-stationary point to (VOP).
@@
The Assumption 3.1(i) is a standard condition for nonconvex cases. Moreover, the Assumption 3.1(ii) holds for many descent methods in vector (multiobjective) optimization, such as $K$-steepest descent method, Barzilai-Bowein descent method [7] and Newton method.
@@
$-\infty<\max_{c^{*}\in C}\left\langle{c^{*},F^{*}-F(x^{0})}\right\rangle\leq\max_{c^{*}\in C}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in C_{k}}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\sum_{k=0}^{\infty}\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle.$
This, together with the fact that $G_{k}(x^{k+1})\preceq_{K}G_{k}(x^{k})=0$, implies $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$. For the accumulation point $x^{*}$, there exists an infinite index set $\mathcal{K}$ such that $x^{k}\stackrel{{\scriptstyle\mathcal{K}}}{{\longrightarrow}}x^{*}$. Therefore, by Assumption 3.1(ii), we conclude that $x^{*}$ is a $K$-stationary point. ∎
@@ 3.2.2 Strong convergence
Suppose that Assumption 3.1 holds and $F(\cdot)$ is $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$. Then, the following statements hold:
@@
    $u_{0}(x^{k})\leq\frac{\ell_{\max}R^{2}}{2k},~{}\forall k\geq 1$, where $\ell_{\max}:=\max_{c^{*}\in C_{k}}\left\langle{c^{*},\bm{\ell}}\right\rangle$, $R:=\{\left\lVert{x-y}\right\rVert:x,y\in\mathcal{L}_{F}(x^{0})\}$, and
    $u_{0}(x^{k}):=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$
@@
Furthermore, we use the fact that $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$ to get
@@
(ii) Since $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$, we use inequality (5) and the definition of $\ell_{\max}$ to obtain
@@
Denote $\hat{c}^{*}_{k}:={\sum_{s=0}^{k-1}c^{*}_{s}}/{k}$, it follows by the convexity of $C_{k}$ that $\hat{c}^{*}_{k}\in C_{k}$. Therefore, we conclude that
@@
Select $y^{k}\in\mathop{\arg\max}_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$, it holds that
$u_{0}(x^{k})$ $=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle=\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(y^{k})}\right\rangle$
@@ 3.2.3 Linear convergence
Suppose that Assumption 3.1(ii) holds $F(\cdot)$ is strongly $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\prec_{K}\bm{\mu}$. Then, the following statements hold:
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{\max\limits_{c^{*}\in C_{k}}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
@@
4 Descent methods for VOPs without line search
@@ 4.1 $K$-steepest descent method for VOPs without line search
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C_{k}}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}.$ (6)
Select a vector $e\in{\rm int}(K)$, and denote $C_{e}=\{c^{*}\in K^{*}:\left\langle{c^{*},e}\right\rangle=1\}$. If we set $C_{k}=C_{e}$ in (6), then the $K$-steepest descent direction can be reformulated as the optimal solution of
@@
Since $F(\cdot)$ is strongly $K$-convex, it follows that $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$ and Assumption 3.1 holds in this case. By setting $C=C_{e}$, Theorem 3.3 (i) and (ii) reduce to the assertions (i) and (ii), respectively. ∎
@@
If $K=\mathbb{R}^{m}_{+}$ and $e=(1,...,1)^{T}$, then $C_{e}$ is a base of +$\mathbb{R}^{m}_{+}$, the convergence rate in Lemma 4.1 reduces to that of [44, Theorem 5.3] with $g(\cdot)=0$. It is worth noting that the rate of convergence is related to $e$, to the best of our knowledge, apart from $e=(1,...,1)^{T}$, it remains an open problem for the better choice of the parameter in MOPs.
@@ 4.3 Equivalence between Algorithms 2 and 3
Similar to Lemma 4.2, Algorithm 4 enjoys the following improved linear convergence.
@@
If $K=\mathbb{R}^{m}_{+}$, and $\Delta_{m}^{\bm{\ell}}:=\{c^{*}\in\mathbb{R}^{m}_{++}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}$, the Algorithm 4 reduces to [8, Algorithm 5] with $g(\cdot)=0$. Interestingly, the relations between Algorithms 2, 3 and 4 only depends on the choice of $e$. If $e=\bm{\ell}$, Algorithms 2, 3 and 4 are equivalent. As a result, for the open problem of the better choice of $e$ in Remark 4.2, we give a partially positive answer with $e=\bm{\ell}$.
@@
Although Algorithms 3 and 4 both exhibit similar improved linear convergence, the computational cost of solving subproblems is generally lower in Algorithm 4. Details on solving these subproblems will be provided in Section 6.
5 Descent methods for VOPs with line search
In the previous section, the majorization-minimization optimization methods were developed using majorizing surrogate functions; however, these surrogates may be overly conservative due to reliance on global upper bounds. From the perspective of majorization-minimization, selecting a non-majorizing surrogate function could potentially enhance performance.
@@ 7             $x^{k+1}:=x^{k}+t_{k}d^{k}$
$G_{k,e/t_{\min}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2t_{\min}}\left\lVert{x-x^{k}}\right\rVert^{2}e.$ (11)
We can show that $G_{k,e/t_{\min}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
@@
Let $G_{k,e/t_{\min}}(\cdot)$ be defined as (11). Then, the following statements hold.
@@
    If $F(\cdot)$ is $K$-convex, then $G_{k,e/t_{\min}}(\cdot)\in\mathcal{S}_{e/t_{\min},e/t_{\min}}(F,x^{k})$.
@@
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,e/t_{\min}}(\cdot)\in\mathcal{S}_{e/t_{\min}-\bm{\mu},e/t_{\min}}(F,x^{k})$.
Proof.
@@
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$, the convergence rate in Lemma 5.1(ii) reduces to that of [19, Theorem 4.2].
@@ 8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
It is worth noting that we don’t specify how to select $e_{k}$ in Algorithm 6. This naturally raises the question: what role does $e_{k}$ play in determining the convergence rate?
Firstly, we derive the lower bound of stepsize in each iteration.
@@
The stepsize generated in Algorithm 6 satisfies $t_{k}\geq t_{k}^{\min}:=\min\left\{\min\limits_{c^{*}\in C_{e}}\frac{\gamma\left\langle{c^{*},e_{k}}\right\rangle}{\left\langle{c^{*},\bm{\ell}}\right\rangle},1\right\}$.
Proof.
The result can be obtained by using the similar arguments as in the proof of Proposition 5.1. ∎
@@
$G_{k,e_{k}/t_{k}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2t^{\min}_{k}}\left\lVert{x-x^{k}}\right\rVert^{2}e_{k}.$ (12)
We can show that $G_{k,e_{k}/t^{\min}_{k}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
@@
Let $G_{k,e_{k}/t^{\min}_{k}}(\cdot)$ be defined as (12). Then, the following statements hold.
@@
    If $F(\cdot)$ is $K$-convex, then $G_{k,e_{k}/t^{\min}_{k}}(\cdot)\in\mathcal{S}_{e_{k}/t^{\min}_{k},e_{k}/t^{\min}_{k}}(F,x^{k})$.
@@
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,e_{k}/t^{\min}_{k}}(\cdot)\in\mathcal{S}_{e_{k}/t^{\min}_{k}-\bm{\mu},e_{k}/t^{\min}_{k}}(F,x^{k})$.
@@
The following results show that $e_{k}$ plays a significant role in the convergence rate of Algorithm 6.
@@
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t^{\min}_{k}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
  • (iii)
    If $e_{k}=e$, we have
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-t_{\min}{\mu_{\min}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0,$
    where $\mu_{\min}:=\min_{c^{*}\in C_{e}}\left\langle{c^{*},\bm{\mu}}\right\rangle$.
  • (iv)
    For any $e_{k}\in\mathrm{int}K$, we have
    $\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t^{\min}_{k}}\right\rangle}\leq\frac{\gamma}{\kappa_{F,\preceq_{K}}}.$ (13)
    Moreover, the equality holds with $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$. In these cases, we have
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-{\gamma}/{\kappa_{F,\preceq_{K}}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0.$ (14)
@@
The assertions (i) and (ii) can be obtained by using the similar arguments as in the proof of Lemma 4.1. By substituting $e_{k}=e$ into (ii), we can obtain the assertion (iii). Next, we prove the assertion (iv). Since $C_{e}$ is a compact set, there exists a vector $c_{0}^{*}\in C_{e}$ such that $1/\kappa_{F,\preceq_{K}}={\left\langle{c_{0}^{*},\bm{\mu}}\right\rangle}/{\left\langle{c_{0}^{*},\bm{\ell}}\right\rangle}.$ On the other hand, by the definition of $t_{k}^{\min}$ we can deduce
$\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t^{\min}_{k}}\right\rangle}\leq\frac{\left\langle{c_{0}^{*},\bm{\mu}}\right\rangle}{\left\langle{c_{0}^{*},e_{k}}\right\rangle}\frac{\gamma\left\langle{c_{0}^{*},e_{k}}\right\rangle}{\left\langle{c_{0}^{*},\bm{\ell}}\right\rangle}=\frac{\gamma}{\kappa_{F,\preceq_{K}}}.$
Then the relation (13) follows. The equality can be obtain by substituting $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$ into the left-hand side of (13) . Moreover, The equality leads to the relation (14). ∎
@@
As described in Lemma 5.2(iv), by setting $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$, we can derive the optimal linear convergence rate for Algorithm 6. Moreover, the convergence rate reduces to that of Lemma 4.3(ii) with constant $\gamma$. Intuitively, to explore the local curvature information of $F(\cdot)$, we can devise a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$ with $\bm{\mu}\preceq_{K}e_{k}\preceq_{K}\bm{\ell}$. In this case, the linear convergence rate of Algorithm 6 can be further improved by using a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$.
@@
6 Descent methods for VOPs with polyhedral cones